OpenStudy (anonymous):

The number of people in a town of 10,000 who have heard a rumor started by a small group of people is given by the following function : N(t) = 10,000 / 5 + 1245e^-0.97t How long will it be until 1,000 people in the town have heard the rumor?

OpenStudy (anonymous):

oh this one again

OpenStudy (anonymous):

you want a long method, or go straight to the answer?

OpenStudy (anonymous):

i had the decimal in the wrong place, is this i think\[\huge N(t)=\frac{10,000}{5+1245e^{-0.97t}}\]

OpenStudy (anonymous):

Yeah, I solved it but the answer seemed wrong. The answer would be fine and I can just use what I know to show the work!

OpenStudy (anonymous):

you set it equal to 1000 and start with \[\frac{10,000}{5+1245e^{-0.97t}}=1000\]

OpenStudy (anonymous):

That's what I did, but I got a decimal for my answer although my math could be wrong ha

OpenStudy (anonymous):

then some algebra \[10,000=1000(5+1245e^{-.97t})\\ 10=5+e^{-.97t}\\ 5=e^{-.97t}\]

OpenStudy (anonymous):

the in log form \[\ln(5)=-.97t\] so \[t=\frac{\ln(5)}{-.97}\]

OpenStudy (anonymous):

ok i made a mistake somewhere

OpenStudy (anonymous):

ooho i see it

OpenStudy (anonymous):

\[10,000=1000(5+1245e^{-.97t})\\ 10=5+1245e^{-.97t}\\ 5=1245e^{-.97t}\]

OpenStudy (anonymous):

so \[\frac{5}{1245}=e^{-.97t}\] and then \[\frac{1}{249}=e^{-.79t}\] \[\ln(\frac{1}{249}=-.97t\\ t=\frac{\ln(1/249)}{-.97}\]

OpenStudy (anonymous):

i get about \(5.7\)

OpenStudy (anonymous):

@satellite73 Thank you so much! My computer was acting up so I had to restart it. But I really appreciate your help!